60 likes | 204 Views
Activities Status and Plan for the week. D. Giugni. Plan of the week. Plan of the week. WELDING Last batch is at CERN!!! HeX Pipe pressure/leak testing should be finished today (Mito) Michal should be back today. Next HeX’s set is A56 (then C78 ) (Michal)
E N D
Activities Status and Plan for the week D. Giugni Monday Meeting 2012-09-3
Plan of the week • WELDING • Last batch is at CERN!!! • HeX • Pipe pressure/leak testing should be finished today (Mito) • Michal should be back today. Next HeX’s set is A56 (then C78) (Michal) • Bellows are missing! Needs to produce at 2 sets (Mito + Jan) • Urgently needs to swage C56 and apply extension pipes to C56 + A78 (Michal) • Comments: Michal, Mito
Plan of the week • Final assembly • A12 and C12 are completed but still need to be surveyed (see later) • C56 needs to wrap up LEMO and ER_Bundles Hans, Mito • A78: • Panel OSP A8 has been repaired… needs to be tested (Daniel) • Planned to go all the way through… (Valery) • Panel OSP A6 needs to replace the D-Bundle with a B-Bundle (Valery) • Comments: Valery, Hans, Maurice, Bryan and Daniel
SURVEY • Problems with the Romer arm. Tool got aligned only last Friday • This week we plan to survey all the available backbones and verify whether the fix works. • Fix is then applied to all the backbones . • A12, C12, C56 and A78 are surveyed afterword and eventually fixed. Comments: Hans, Maurice Monday Meeting 2012-09-3
DBM Bundles and NTC Sensors • Prototype Power+Signal Bundles have been soldered to the DBM PP0 board. Found a swap in the connectivity Corrected! • ~10 NTC pigtails have been produced. THIS WEEK: Plan is not crystal clear: few options: • Produce the second DBM “production” bundle … but LEMO AWG36 pins are still missing… • Work out a routing solution for the NTC on any of the available nSQP and define the final lengths of the NTCpigtails and the connector position. • Move Olga to help Mito in the Pipes Lab. Monday Meeting 2012-09-3